# Learning topological states from randomized measurements using variational tensor network tomography

### v3

@@ B.1 Matrix product states
[…]
@@
[…]
@@
where $[\psi_{\chi}(A)]$ denotes the ray representing the equivalence class of the state vector $\psi_{\chi}(A)$ under rescaling by nonzero complex numbers.
[…]
@@ B.2 Differential geometry
Let us begin by briefly recalling the concept of an $r$-dimensional differentiable manifold $M$. We use $\theta\in M$ to denote elements on the manifold, where, in the subsequent sections, $\theta$ parametrizes some underlying probability distribution. Formally, any local neighbourhood $M_{i}\subset M$ can be related to Euclidean space $\mathbb{R}^{r}$. This relationship is established through an injective map $\varphi_{i}:M_{i}\rightarrow\mathbb{R}^{r}$, known as a chart. A chart provides a local coordinate system near each point in its neighbourhood. Intuitively, the manifold is covered by a collection of neighbourhoods $M_{i}$ such that the union $\bigcup_{i\in I}M_{i}=M$ represents the entire manifold.
The collection of charts, known as an atlas, allows us to study functions defined on $M$ using the usual machinery from calculus defined on $\mathbb{R}^{r}$. Moreover, for $M$ to be called differentiable, any composition $\varphi_{i}^{-1}\circ\varphi_{j}$ defined on any two overlapping patches $M_{i}\cap M_{j}\neq\emptyset$ must be differentiable. This allows us to define the tangent vectors $X$ at any point $\theta\in M$. In alignment with Ref. [37], we adopt a geometric perspective of tangent vectors by representing them as column vectors $(X_{1},\cdots,X_{r})^{\top}$. The components of the vector $X$ depend on the local coordinates in $\mathbb{R}^{r}$, defined by the chosen chart $\varphi$ [^3] Tangent vectors to $M$ at points $\theta$ can be naturally seen as directional derivatives. For a smooth function $f:M\rightarrow\mathbb{R}$ defined on the manifold, if $X$ is a tangent vector to the manifold $M$ at $\theta$, then the directional derivative (defined in Def. 6) is given by $D_{X}(f)=X(f)$. For what follows, it is convenient to directly refer to tangent vectors. The space of all tangent vectors at $\theta$ is called the tangent space $\mathrm{T}_{\theta}M$ and it is isomorphic to $\mathbb{R}^{r}$. Consequently, each tangent vector can be expressed as $X=\sum_{a=1}^{r}X_{a}e_{a}$ where $\{e_{a}\}$ is a basis of $\mathrm{T}_{\theta}M$.
@@
In the discussion that follows, we will use $g$ to denote the metric $g_{\theta}$, as it is understood that the metric is defined for the tangent space at each point $\theta\in M$. Additionally, we adopt the notation |$_{g}$ to represent the inner product of any two tangent vectors, which is computed according to the metric $g$.
@@
for all $X,Y,Z\in\Gamma(\mathrm{T}M)$ and all smooth functions $f,h:M\rightarrow\mathbb{R}$ defined on the manifold. [,][$\bullet$,$\bullet$][ ∙ , ∙ ] is the Lie bracket defined by $[X,Y](f):=\mathrm{D}_{X}\mathrm{D}_{Y}f-\mathrm{D}_{Y}\mathrm{D}_{X}f$.
@@
A special case of parallel vector fields is given by geodesic curves. The velocity vector $\dot{\gamma}$ of a geodesic curve $\gamma:I\rightarrow M$ (with $I\subseteq\mathbb{R}$ some parameter interval) is parallel along $\gamma$ itself. For simplicity, we will take the unit interval $I=[0,1]$ because a curve with any other interval can always be reparameterized. The notion of a geodesic extends the concept of a straight line from Euclidean space to curved spaces like manifolds. In general, there can be multiple geodesic curves connecting two points on a manifold [^4] For example, the great circle on the sphere $S^{2}$ is a geodesic curve. Two points on the great circle separate the circle into two arcs. The shorter arc is the shortest connecting geodesic, and the longer arc is the longer connecting geodesic.. For our purpose, we only consider the curves that represent the shortest path between two points on the manifold, parametrized by arc length on the unit interval.
@@
[…]
@@
00 0 $=\nabla_{\dot{\gamma}}\dot{\gamma}$ (27)
@@
[…]
@@ B.3 Projective Hilbert space as a manifold
Note that $\mathsf{P}(\mathcal{H})$ can be naturally seen as a complex manifold of complex dimension $12^{n}-$ 1, although it could, at the same time, also be defined as a real manifold of dimension $2(2^{n}-1)$. Following Ref. [81] (Sec. 8.2) and Ref. [37] (Sec. IIA), we can take the tangent space of the complex manifold $\mathsf{P}(\mathcal{H})$ at some point $[\psi]\in\mathsf{P}(\mathcal{H})$ as the real vector space $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$ of dimension $2(2^{n}-1)$. To avoid confusion, we remark that it can be useful to consider the complexification $\mathrm{T}^{\mathbb{C}}_{[\psi]}\mathsf{P}(\mathcal{H})$ (i.e., the complex span of any basis of $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$) of this vector space to define a set of so-called holomorphic tangent vectors. Note, however, that $\mathrm{T}^{\mathbb{C}}_{[\psi]}\mathsf{P}(\mathcal{H})$ has complex dimension $2(2^{n}-1)$ and is therefore over-parametrized.
For our result, we only rely on the manifold $\mathsf{P}(\mathcal{H})$ having the structure of a Riemannian manifold and therefore, we will only work with $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$. However, as we will see, $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$ is closely related to the Hilbert space $\mathcal{H}$ itself. Hence, it is more natural to view the $2(2^{n}-1)$-dimensional real tangent vectors as $12^{n}-$ 1-dimensional complex vectors. Those correspond to vectors in $\mathcal{H}$ that are orthogonal to $[\psi]$. We stress that these vectors are not elements of the complexified tangent space (which would be complex vectors of dimension $2(2^{n}-1)$).
@@
[…]
@@
Let $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ be two tangent vectors in $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$ and $|\psi\rangle$ be the normalized representation of $[\psi]$. The Fubini-Study metric is defined as
@@
[…]
@@
Note that this Fubini-Study distance is known to be related to the angle between the two states $[\phi]$ and $[\psi]$. This is often stated without any proof, but rather the intuitive motivation is that the shortest path in projective space corresponds to the arc length of the enclosed segment of the unit circle containing $[\phi]$ and $[\psi]$. Since we define the state distance via the curve length of the shortest connecting geodesic, we state an explicit expression of the Fubini-Study distance below.
@@ B.4 Manifold of matrix product states
[…]
@@ B.5 Maximum likelihood estimation on manifold
$\mathcal{L}^{D}_{\text{NLL}}:M\leavevmode\nobreak$ $\longrightarrow\mathbb{R},$ (42)
@@
    as $\delta\rightarrow 0$, where $B_{\theta_{0}}(\delta)$ is the set of all points with distance at most $\delta$ from $\theta_{0}$ (the “$\delta$-Ball” around $\theta_{0}$).
@@
[…]
@@ B.6 Main result: A concentration inequality for infidelity
[…]
@@
$\mathcal{L}^{D}_{\text{NLL}}:\tilde{M}\leavevmode\nobreak$ $\longrightarrow\mathbb{R},$ (53)
@@
[…]
$[{\psi}^{D}]=\text{argmin}_{[\psi]\in\tilde{M}}\mathcal{L}^{D}_{\text{NLL}}\left([\psi]\right).$ (55)
Using the notation established in Appendix B.5, we use $[\hat{\psi}^{(N)}]$ to denote the random variable of the MLE estimator, which is the minimizer of the $N$-sample NLL for a dataset random variable ${D}^{(N)}$. Using Fact 2 we find a concrete expression of the estimating form by taking covariant derivative of the NLL as follows.
@@
Let ${D}^{(N)}=\{(U_{i},b_{i})\}_{i=1}^{N}$ be a sequence of i.i.d. samples from $\mathcal{D}$ defined by the target state $[\phi]$ and a unitary ensemble $\mathcal{U}$, tomographically complete in the space of variational states on $\tilde{M}$. There exists a sequence of $N$-sample NLL minimizers $([\hat{\psi}^{(N)}])_{N=1}^{\infty}$ such that $\sum_{i=1}^{N}\omega((U_{i},b_{i}),[\hat{\psi}^{(N)}])=0$ for all $N$ and the estimator converges to the target state in probability $[\hat{\psi}^{(N)}]\overset{\mathrm{p.}}{\longrightarrow}[\phi]$.
@@
In the infinite-sample limit, he NLL is given by
@@
where $|\phi\rangle$ and $|\psi\rangle$ are normalized representations of their rays. The NLL loss function achieves the minimum denoted by $[\hat{\psi}]$  [^6] The minimum is given by the Shannon entropy of the target distribution $\mathcal{D}$, calculated as $-\sum_{(U,b)}\mathcal{D}[(U,b),[\phi]]\log\mathcal{D}[(U,b),[\phi]]$. This follows from Gibbs’ inequality, which, for two discrete probability distributions $\{p_{1},\cdots,p_{N}\}$ and $\{q_{1},\cdots,q_{N}\}$, states that $-\sum_{i=1}^{N}p_{i}\log p_{i}\leq-\sum_{i=1}^{N}p_{i}\log q_{i}$. The equality is achieved when two distributions are the same $p_{i}=q_{i}$. if the probabilities agree for all measurement outcomes
@@
[…]
@@
The expected Hessian at the target state $[\phi]$ is invertible and finite.
@@
The Hessian at $[\hat{\psi}^{(N)}]$ in a small neighbourhood near the target converges uniformly to the Hessian at $[\phi]$ in expectation with respect to $\mathcal{D}$.
@@
Let $\mathcal{D}$ be the distribution corresponding to the target state $[\phi]$ (defined in Eq. (52)), and ${D}^{(N)}=\{(U_{i},b_{i})\}_{i=1}^{N}$ be a sequence of i.i.d. samples from $\mathcal{D}$. Let $[\hat{\psi}^{(N)}]$ be the minimizer of the negative log-likelihood function and $\omega$ its assoicated estimating form as defined in Corollary 1. Define $\gamma$ as the shortest connecting geodesic such that $\gamma(0)=[\phi]$, a point $\gamma(1)=[\hat{\psi}^{(N)}]$, and $\dot{\gamma}$ as the velocity vector along the curve. Let the velocity vector be decomposed in the parallel transporte—along $\gamma$—basis vectors $\{e_{a}(t)\}$ as $\dot{\gamma}=\sum_{a}\Delta_{a}e_{a}(t)$. Suppose the conditions of Proposition 1, Assumption 1, and Assumption 2 are satisfied. By Lemma 3, the velocity vector’s components $\{\Delta_{a}\}_{a=1}^{r}$ converge in distribution to a multivariate normal distribution
@@
[…]
@@
Given two states $[\psi],[\phi]$ in $\tilde{M}$, the projective space of MPS (see Lemma 2), denote the shortest connecting geodesic by $\gamma:[0,1]\rightarrow\tilde{M}$, such that $\gamma(0)=[\psi]$ and $\gamma(1)=[\phi]$, as defined in Def. 8. The infidelity of $[\psi]$ and $[\phi]$ is bounded by the length of the geodesic:
@@
where |$_{\tilde{g}}$ denotes the induced Fubini-Study metric on $\tilde{M}$.
@@
[…]
@@
Let $\tilde{M}$ be the manifold of full-rank MPS of a maximum bond dimensions $\chi_{\max}$, as per Eq. (40), and let $[\phi]\in\tilde{M}$ be some target state. Let $D^{(N)}=\{(U_{i},\,b_{i})\}_{i=1}^{N}$ be the random variable representing a dataset of $N$ samples drawn from the distribution $\mathcal{D}[(U,b),[\phi]]$ defined in Eq. (52). Denote by $[\hat{\psi}^{(N)}]$ the random variable representing the minimizer, in $\tilde{M}$, of the $N$-sample negative log-likelihood defined by $D^{(N)}$, and let $\omega$ be its associated estimating form as defined in Corollary 1. Assume the conditions of Proposition 1, Assumption 1, and Assumption 2 are satisfied. Then, for any $\delta\in(0,1]$, there exists $N_{0}$ such that for all $N>N_{0}$, one has

### v4

@@ B.1 Matrix product states
The set $\{\mbox{$|b\rangle$}\}$ represents the computational basis, where the bitstring $b\in\{0,1\}^{n}$, unless specified otherwise. Note that in the main text, we also use the notation $|\psi(A)\rangle$ to refer to the physical state, which is inferred from the context.
@@
which contains all rays represented by the equivalence classes $\mbox{$|\psi\rangle$}]$. Two nonzero vectors $\mbox{$|\psi\rangle$},\mbox{$|\phi\rangle$}\in\mathcal{H}$ are equivalent if and only if $\mbox{$|\psi\rangle$}=\lambda\mbox{$|\phi\rangle$}$ for some $\lambda\in\operatorname{GL}(1,\mathbb{C})$, the group of nonzero complex numbers. For simplicity, we use the notation $\psi]$ to denote $\mbox{$|\psi\rangle$}]$.
@@
where $\psi_{\chi}(A)]$ denotes the ray representing the equivalence class of the state vector $\psi_{\chi}(A)$ under rescaling by nonzero complex numbers.
Throughout the rest of this section, what we refer to by $|\psi\rangle$ or $\psi]$ should be understood from the context as denoting either a normalized state vector or a ray in the projective Hilbert space, in distinction to the projective mapping $\tilde{\psi}_{\chi}$. Note that the image $\tilde{\psi}_{\chi}(\mathbb{A}_{\chi})$ represents only a subspace of $\mathsf{P}(\mathcal{H})$ and does not possess the structure of a vector space  [^2] For instance, consider the sum of two MPS of bond dimension $\chi$ as $({\mbox{$|\psi_{\chi}\rangle$}+\mbox{$|\phi_{\chi}\rangle$}})/{\sqrt{2}}$. This new state has bond dimension up to $2\chi$, which is beyond the original space defined for bond dimension $\chi$.. In the following sections, we will provide constraints on the domain of MPS tensors such that their image under $\tilde{\psi}_{\chi}$ (and $\psi_{\chi}$) forms a differentiable manifold within $\mathsf{P}(\mathcal{H})$. Equipping the set of MPS with such an additional structure, we will be able to rigorously analyze the accuracy of the MLE solution on the space of a manifold, as detailed in Appendix B.5.
@@ B.2 Differential geometry
Let us begin by briefly recalling the concept of an $r$-dimensional differentiable manifold $M$. We use $\theta\in M$ to denote elements on the manifold, where, in the subsequent sections, $\theta$ parametrizes some underlying probability distribution. Formally, any local neighborhood $M_{i}\subset M$ can be related to Euclidean space $\mathbb{R}^{r}$. This relationship is established through an injective map $\varphi_{i}:M_{i}\rightarrow\mathbb{R}^{r}$, known as a chart. A chart provides a local coordinate system near each point in its neighborhood. Intuitively, the manifold is covered by a collection of neighborhoods $M_{i}$ such that the union $\bigcup_{i\in I}M_{i}=M$ represents the entire manifold.
The collection of charts, known as an atlas, allows us to study functions defined on $M$ using the usual machinery from calculus defined on $\mathbb{R}^{r}$. Moreover, for $M$ to be called differentiable, any composition $\varphi_{i}^{-1}\circ\varphi_{j}$ defined on any two overlapping patches $M_{i}\cap M_{j}\neq\emptyset$ must be differentiable. This allows us to define the tangent vectors $X$ at any point $\theta\in M$. In alignment with Ref. [37], we adopt a geometric perspective of tangent vectors by representing them as column vectors $(X_{1},\cdots,X_{r})^{\top}$. The components of the vector $X$ depend on the local coordinates in $\mathbb{R}^{r}$, defined by the chosen chart $\varphi$. [^3] Tangent vectors to $M$ at points $\theta$ can be naturally seen as directional derivatives. For a smooth function $f:M\rightarrow\mathbb{R}$ defined on the manifold, if $X$ is a tangent vector to the manifold $M$ at $\theta$, then the directional derivative (defined in Def. 6) is given by $D_{X}(f)=X(f)$. For what follows, it is convenient to directly refer to tangent vectors. The space of all tangent vectors at $\theta$ is called the tangent space $\mathrm{T}_{\theta}M$ and it is isomorphic to $\mathbb{R}^{r}$. Consequently, each tangent vector can be expressed as $X=\sum_{a=1}^{r}X_{a}e_{a}$ where $\{e_{a}\}$ is a basis of $\mathrm{T}_{\theta}M$.
@@
In the discussion that follows, we will use $g$ to denote the metric $g_{\theta}$, as it is understood that the metric is defined for the tangent space at each point $\theta\in M$. Additionally, we adopt the notation $\mbox{$\langle\bullet|\bullet\rangle$}_{g}$ to represent the inner product of any two tangent vectors, which is computed according to the metric $g$.
@@
for all $X,Y,Z\in\Gamma(\mathrm{T}M)$ and all smooth functions $f,h:M\rightarrow\mathbb{R}$ defined on the manifold. $\bullet$$\bullet$] is the Lie bracket defined by $X,Y](f):=\mathrm{D}_{X}\mathrm{D}_{Y}f-\mathrm{D}_{Y}\mathrm{D}_{X}f$.
@@
A special case of parallel vector fields is given by geodesic curves. The velocity vector $\dot{\gamma}$ of a geodesic curve $\gamma:I\rightarrow M$ (with $I\subseteq\mathbb{R}$ some parameter interval) is parallel along $\gamma$ itself. For simplicity, we will take the unit interval $I=[0,1]$ because a curve with any other interval can always be reparametrized. The notion of a geodesic extends the concept of a straight line from Euclidean space to curved spaces like manifolds. In general, there can be multiple geodesic curves connecting two points on a manifold [^4] For example, the great circle on the sphere $S^{2}$ is a geodesic curve. Two points on the great circle separate the circle into two arcs. The shorter arc is the shortest connecting geodesic, and the longer arc is the longer connecting geodesic.. For our purpose, we only consider the curves that represent the shortest path between two points on the manifold, parametrized by arc length on the unit interval.
@@
$\|\dot{\gamma}(t)\|$ $:=\sqrt{\mbox{$\langle\dot{\gamma}|\dot{\gamma}\rangle$}_{g}}$ (26)
@@
$0$ $=\nabla_{\dot{\gamma}}\dot{\gamma}$ (27)
@@
$=\sqrt{\mbox{$\langle\dot{\gamma}|\dot{\gamma}\rangle$}_{g}},$
@@ B.3 Projective Hilbert space as a manifold
Note that $\mathsf{P}(\mathcal{H})$ can be naturally seen as a complex manifold of complex dimension $2^{n}-1$, although it could, at the same time, also be defined as a real manifold of dimension $2(2^{n}-1)$. Following Ref. [81] (Sec. 8.2) and Ref. [37] (Sec. IIA), we can take the tangent space of the complex manifold $\mathsf{P}(\mathcal{H})$ at some point $\psi]\in\mathsf{P}(\mathcal{H})$ as the real vector space $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$ of dimension $2(2^{n}-1)$. To avoid confusion, we remark that it can be useful to consider the complexification $\mathrm{T}^{\mathbb{C}}_{[\psi]}\mathsf{P}(\mathcal{H})$ (i.e., the complex span of any basis of $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$) of this vector space to define a set of so-called holomorphic tangent vectors. Note, however, that $\mathrm{T}^{\mathbb{C}}_{[\psi]}\mathsf{P}(\mathcal{H})$ has complex dimension $2(2^{n}-1)$ and is therefore over-parametrized.
For our result, we only rely on the manifold $\mathsf{P}(\mathcal{H})$ having the structure of a Riemannian manifold and therefore, we will only work with $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$. However, as we will see, $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$ is closely related to the Hilbert space $\mathcal{H}$ itself. Hence, it is more natural to view the $2(2^{n}-1)$-dimensional real tangent vectors as $2^{n}-1$-dimensional complex vectors. Those correspond to vectors in $\mathcal{H}$ that are orthogonal to $\psi]$. We stress that these vectors are not elements of the complexified tangent space (which would be complex vectors of dimension $2(2^{n}-1)$).
@@
The tangent space $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})=\mathcal{H}/\!\sim$ at a base point denoted by $\psi]$ is the space of all tangent vectors that satisfy the following equivalence relation. Let $\mbox{$|\psi_{1}\rangle$},\mbox{$|\psi_{2}\rangle$}\in\mathcal{H}$ be two state vectors in the Hilbert space,
$\mbox{$|\psi_{1}\rangle$}\sim\mbox{$|\psi_{2}\rangle$}:\!\iff\mbox{$|\psi_{1}\rangle$}-\mbox{$|\psi_{2}\rangle$}=\alpha\mbox{$|\psi\rangle$},$ (30)
@@
Let $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ be two tangent vectors in $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$ and $|\psi\rangle$ be the normalized representation of $\psi]$. The Fubini-Study metric is defined as
@@
$(\mbox{$|{\psi_{1}}\rangle$},\mbox{$|{\psi_{2}}\rangle$})$ $\longmapsto\mathop{\mathrm{Re}}\big{[}\mbox{$\langle\psi_{1}|\psi_{2}\rangle$}-\langle\psi_{1}\lvert\psi\rangle\langle\psi\rvert\psi_{2}\rangle\big{]}.$ (32)
@@
Note that this Fubini-Study distance is known to be related to the angle between the two states $\phi]$ and $\psi]$. This is often stated without any proof, but rather the intuitive motivation is that the shortest path in projective space corresponds to the arc length of the enclosed segment of the unit circle containing $\phi]$ and $\psi]$. Since we define the state distance via the curve length of the shortest connecting geodesic, we state an explicit expression of the Fubini-Study distance below.
@@ B.4 Manifold of matrix product states
Note that in Ref. [37], it is demonstrated that $\tilde{M}$ is actually a Kähler manifold when one endows it with the complex-valued Fubini-Study metric, which is shown to be a Hermitian metric. This introduces additional structure—studied in complex geometry—to the manifold. Since we only rely on the variational set being a well-defined Riemannian manifold, we ignore the additional structure here and only define the real-valued Fubini-Study metric in alignment with Def. 11 by inducing the Fubini-Study metric from the embedding space $\mathsf{P}(\mathcal{H})$. The projective MPS manifold is a submanifold $\tilde{M}\subset\mathsf{P}(\mathcal{H})$ of the projective Hilbert space, which is equipped with the Fubini-Study metric defined in Def. 11. For each point $\psi]\in\tilde{M}$, the associated tangent space $\mathrm{T}_{[\psi]}\tilde{M}$ is a subspace of $\mathrm{T}_{[\psi]}\mathsf{P}(\mathcal{H})$. As a result, one can define the induced Fubini-Study metric $\tilde{g}$ on $\tilde{M}$ such that for any $\psi]\in\tilde{M}$ and any $\mbox{$|\phi_{1}\rangle$},\mbox{$|\phi_{2}\rangle$}\in\mathrm{T}_{[\psi]}\tilde{M}$, one has
$\mbox{$\langle\phi_{1}|\phi_{2}\rangle$}_{\tilde{g}}=\mbox{$\langle\phi_{1}|\phi_{2}\rangle$}_{g},$ (41)
@@ B.5 Maximum likelihood estimation on manifold
$\mathcal{L}^{D}_{\text{NLL}}:M\leavevmode\nobreak\$$\longrightarrow\mathbb{R},$ (42)
@@
    as $\delta\rightarrow 0$, where $B_{\theta_{0}}(\delta)$ is the set of all points with distance at most $\delta$ from $\theta_{0}$ (the $\delta$-Ball” around $\theta_{0}$).
@@
where $\Sigma=(\mathrm{H}^{\dagger})^{-1}\Gamma\mathrm{H}$ and the Gram matrix $\Gamma_{a,b}=\mathbb{E}_{x\sim\mathcal{P}_{\theta_{0}}}\big{[}\mbox{$\langle\omega(x,\theta_{0})|e_{a}\rangle$}\mbox{$\langle\omega(x,\theta_{0})|e_{b}\rangle$}\big{]}$.
@@ B.6 Main result: A concentration inequality for infidelity
Consider the target state ${\phi}]\in\tilde{M}$, with a normalized representative $\mbox{$|\phi\rangle$}\in\mathcal{H}$. Given a unitary ensemble $\mathcal{U}$, the sample space $\mathcal{S}$ consists of all unitary-bitstring pairs $\mathcal{S}=\mathcal{U}\times\{0,1\}^{n}$. Let $\mathcal{D}$ denote the data probability distribution corresponding to the target state $\phi$}] given by
@@
$\mathcal{L}^{D}_{\text{NLL}}:\tilde{M}\leavevmode\nobreak\$$\longrightarrow\mathbb{R},$ (53)
@@
Notice that by introducing a scaling factor $c>1$, the loss associated with any vector can be arbitrarily reduced through the transformation $\mbox{$|\psi\rangle$}\mapsto c\mbox{$|\psi\rangle$}$. However, such rescaling does not affect physical observables and is effectively managed by constraining the analysis to the projective Hilbert space. This highlights the practical value of defining the projective MPS and its associated projective manifold. The MLE estimate is given by the minimizer of the NLL
${\psi}^{D}]=\text{argmin}_{[\psi]\in\tilde{M}}\mathcal{L}^{D}_{\text{NLL}}\left([\psi]\right).$ (55)
Using the notation established in Appendix B.5, we use $\hat{\psi}^{(N)}]$ to denote the random variable of the MLE estimator, which is the minimizer of the $N$-sample NLL for a dataset random variable ${D}^{(N)}$. Using Fact 2 we find a concrete expression of the estimating form by taking covariant derivative of the NLL as follows.
@@
Let ${D}^{(N)}=\{(U_{i},b_{i})\}_{i=1}^{N}$ be a sequence of i.i.d. samples from $\mathcal{D}$ defined by the target state $\phi]$ and a unitary ensemble $\mathcal{U}$, tomographically complete in the space of variational states on $\tilde{M}$. There exists a sequence of $N$-sample NLL minimizers $([\hat{\psi}^{(N)}])_{N=1}^{\infty}$ such that $\sum_{i=1}^{N}\omega((U_{i},b_{i}),[\hat{\psi}^{(N)}])=0$ for all $N$ and the estimator converges to the target state in probability $\hat{\psi}^{(N)}]\overset{\mathrm{p.}}{\longrightarrow}[\phi]$.
@@
In the infinite-sample limit, the NLL is given by
@@
where $|\phi\rangle$ and $|\psi\rangle$ are normalized representations of their rays. The NLL loss function achieves the minimum denoted by $\hat{\psi}$]  [^6] The minimum is given by the Shannon entropy of the target distribution $\mathcal{D}$, calculated as $\sum_{(U,b)}\mathcal{D}[(U,b),[\phi]]\log\mathcal{D}[(U,b),[\phi]]$. This follows from Gibbs’ inequality, which, for two discrete probability distributions $\{p_{1},\cdots,p_{N}\}$ and $\{q_{1},\cdots,q_{N}\}$, states that $\sum_{i=1}^{N}p_{i}\log p_{i}\leq-\sum_{i=1}^{N}p_{i}\log q_{i}$. The equality is achieved when two distributions are the same $p_{i}=q_{i}$. if the probabilities agree for all measurement outcomes
@@
Since the unitary ensemble $\mathcal{U}$ is tomographically complete, Eq. (57) implies $\mbox{$|\hat{\psi}\rangle$}=c\mbox{$|\phi\rangle$}$ up to a global phase $c$, so that $\hat{\psi}]=[\phi]$. ∎
@@
The expected Hessian at the target state $\phi]$ is invertible and finite.
@@
The Hessian at $\hat{\psi}^{(N)}]$ in a small neighborhood near the target converges uniformly to the Hessian at $\phi]$ in expectation with respect to $\mathcal{D}$.
@@
Let $\mathcal{D}$ be the distribution corresponding to the target state $\phi]$ (defined in Eq. (52)), and ${D}^{(N)}=\{(U_{i},b_{i})\}_{i=1}^{N}$ be a sequence of i.i.d. samples from $\mathcal{D}$. Let $\hat{\psi}^{(N)}]$ be the minimizer of the negative log-likelihood function and $\omega$ its assoicated estimating form as defined in Corollary 1. Define $\gamma$ as the shortest connecting geodesic such that $\gamma(0)=[\phi]$, a point $\gamma(1)=[\hat{\psi}^{(N)}]$, and $\dot{\gamma}$ as the velocity vector along the curve. Let the velocity vector be decomposed in the parallel transport—along $\gamma$—basis vectors $\{e_{a}(t)\}$ as $\dot{\gamma}=\sum_{a}\Delta_{a}e_{a}(t)$. Suppose the conditions of Proposition 1, Assumption 1, and Assumption 2 are satisfied. By Lemma 3, the velocity vector’s components $\{\Delta_{a}\}_{a=1}^{r}$ converge in distribution to a multivariate normal distribution
@@
where $\Sigma=(\mathrm{H}^{\dagger})^{-1}\Gamma\mathrm{H}$ and the Gram matrix $\Gamma_{a,b}=\mathbb{E}_{(U,b)\sim\mathcal{D}}\big{[}\mbox{$\langle\omega((U,b),[\phi])|e_{a}\rangle$}\mbox{$\langle\omega((U,b),[\phi])|e_{b}\rangle$}\big{]}$.
@@
Given two states $\psi],[\phi]$ in $\tilde{M}$, the projective space of MPS (see Lemma 2), denote the shortest connecting geodesic by $\gamma:[0,1]\rightarrow\tilde{M}$, such that $\gamma(0)=[\psi]$ and $\gamma(1)=[\phi]$, as defined in Def. 8. The infidelity of $\psi]$ and $\phi]$ is bounded by the length of the geodesic:
@@
where $\mbox{$\langle\bullet|\bullet\rangle$}_{\tilde{g}}$ denotes the induced Fubini-Study metric on $\tilde{M}$.
@@
where $\mbox{$\langle\bullet|\bullet\rangle$}_{g}$ denotes the Fubini-Study metric on $\mathsf{P}(\mathcal{H})$. Note that the choice of the parameter interval $[0,1]$ is without loss of generality because we can reparametrize any connecting curve to an arbitrary closed interval of our choice. Since $\tilde{M}$ is a submanifold of $\mathsf{P}(\mathcal{H})$, we can lift the shortest connecting geodesic $\gamma$ from the projective MPS manifold $\tilde{M}$ into projective Hilbert space $\mathsf{P}(\mathcal{H})$. To be precise, we can formally distinguish the curve $\gamma$ in the two abstract manifolds $\tilde{M}$ and $\mathsf{P}(\mathcal{H})$. For this, we formally identify each point $\gamma(t)\in\tilde{M}$ with its corresponding point in $\gamma^{\prime}(t)\in\mathsf{P}(\mathcal{H})$. Since $\tilde{g}$ is induced by $g$, the lengths of any tangent vectors are the same in both the submanifold and the embedding manifold $\mbox{$\langle\dot{\gamma}|\dot{\gamma}\rangle$}_{\tilde{g}}=\mbox{$\langle\dot{\gamma}^{\prime}|\dot{\gamma}^{\prime}\rangle$}_{g}$ (see Eqs. (103) and (108) in Ref. [37]). Therefore, the lengths of $\gamma$ and $\gamma^{\prime}$ are equal. Choosing the curve to be $\alpha=\gamma^{\prime}$ in Eq. (61), we have proved the result stated in Eq. (59). ∎
@@
Let $\tilde{M}$ be the manifold of full-rank MPS of a maximum bond dimensions $\chi_{\max}$, as per Eq. (40), and let $\phi]\in\tilde{M}$ be some target state. Let $D^{(N)}=\{(U_{i},\,b_{i})\}_{i=1}^{N}$ be the random variable representing a dataset of $N$ samples drawn from the distribution $\mathcal{D}[(U,b),[\phi]]$ defined in Eq. (52). Denote by $\hat{\psi}^{(N)}]$ the random variable representing the minimizer, in $\tilde{M}$, of the $N$-sample negative log-likelihood defined by $D^{(N)}$, and let $\omega$ be its associated estimating form as defined in Corollary 1. Assume the conditions of Proposition 1, Assumption 1, and Assumption 2 are satisfied. Then, for any $\delta\in(0,1]$, there exists $N_{0}$ such that for all $N>N_{0}$, one has